OpenStudy (anonymous):

how many solutions does the equation have? 2(2x+5)=4(x+3) A. one solution B. infinite solutions C. no solution

OpenStudy (campbell_st):

well if you distribute you get 4x + 10 = 4x + 3 do you notice anything... perhaps if I wrote it as y = 4x + 10 and y = 4x + 3

OpenStudy (anonymous):

is it a?

OpenStudy (anonymous):

4 x+10=4 x+12 No Solution.

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

Your welcome.